Find the missing component in each ordered pair using the given equation.
y=2x+1 A.(1,_)
B.(2,_)
C.(_,-3)
D.(_,-5

Answer by rfer(16322)

A)
y=2*1+1
y=2+1
y=3
---------
C)
-3=2x+1
2x+1=-3
subtract 1
2x=-4
x=-2